SpaceRef

SpaceRef


NOAA SATOPS Morning Report: Monday, March 12, 2012

Status Report From: NOAA Office of Satellite Operations
Posted: Monday, March 12, 2012

Customers or users who want to call to ask about problems or report outages can use the following contacts:

- For questions about data quality (noisy data, dropouts, schedules), GOES or POES products, or status of spacecraft can call the SAB Shift Supervisor at 301-763-8444

- To report an outage or document a problem, call the ESPC Operations Crew Lead at 301-457-5218. All LRIT data users are encouraged to report problems to: LRIT@noaa.gov

DOY 072

GOES
GOES-12 DOY70 multiple image frame breaks at 1654Z, 2004Z, and 2154Z due to RFI at WCDA. Under investigation.

All other GOES operations were nominal over the past 72 hours.
No anomalies affecting product delivery via GVAR or s/c health & safety.
For GOES status and other information please see:
http://www.oso.noaa.gov/goes

POES
DOY72 NOAA-19 revs. 15931-15933 were received at ESPC but not at SOCC. Under investigation.

All other POES operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.
For a detailed status and description of POES see:
http://www.oso.noaa.gov/poes

DMSP
All DMSP operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.
For more DMSP Information please see:
http://dmsp.ngdc.noaa.gov/dmsp.html

JASON-2
All JASON-2 operations were nominal over the past 72 hours.
No anomalies affecting product delivery or s/c health & safety.

ESPC Reported data outage and delays
DOY70 Mar 10, 2012 - SSMI data from Ops-54 revs 43303 and 43304 were delayed.

Questions or comments to Ron.Mahmot@noaa.gov

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.